# PsyDI: Towards a Personalized and Progressively In-depth Chatbot for Psychological Measurements

### v1

@@ 3.2 Statement Selection
Following the update of the MBTI Profile, another key challenge emerges in eliciting more representative statement from users through subsequent questioning. Given the preliminary acquisition of user portrait statements, the next process can be divided into two distinct stages: statement selection and multi-turn dialogue with the user. The goal of the former lies in discerning which statements are most indicative of representative responses, thereby delineating the areas requiring in-depth exploration and further analysis. Conversely, the objective of multi-turn dialogue is to engage users, encouraging them to provide more details in the extended question-answers and maintain their interest throughout the assessment.
To determine the most indicative statements, we should exclude those with minimal informational value and those that explicitly describe a specific MBTI type, as these do not warrant further inquiry. Instead, we should focus on statements that are ripe for exploration, particularly those that exhibit ambiguity across multiple MBTI dimensions. For instance, considering a user whose statements could plausibly align with either the INFP or ENFJ types. This ambiguity suggests difficulty in distinguishing between the Introversion/Extraversion (I/E) and Perceiving/Judging (P/J) dimensions. Thus, the subsequent questioning should be designed to delve into the motivations underlying these dimensions. In this context, we design a strategy to select statements that are ambiguous yet highly representative of both the I/E and P/J dimensions, thereby refining our understanding of the user’s MBTI. This process can be formulated as:
$\operatorname*{arg\,min}_{p\in s}\sum_{i=1}^{4}\left(M^{i}_{j}\otimes M^{i}_{k}\right)\cdot\left(\text{rank}_{i}(\mathcal{F}_{\text{MBTI}_{j}}(p))+\text{rank}_{i}(\mathcal{F}_{\text{MBTI}_{k}}(p)\right)$ (5)
The notations $\text{MBTI}_{j}$ and $\text{MBTI}_{k}$ refer to the top-2 MBTI types within the current MBTI profile. $M^{i}_{j}$ denotes the boolean (0/1) value of $\text{MBTI}_{j}$ on the $i$-th MBTI dimension. The symbol $\otimes$ means XOR operation, which is employed to discern the differing dimensions between these two MBTI types. While $\text{rank}_{i}$ signifies the rank of the scores on the $i$-th MBTI dimension across all statements. The objective of Eq. 5 is to pinpoint statements that score highly on all dimensions of ambiguity, enabling subsequent questioning to elucidate the user’s true cognitive preferences.
@@ 3.3 Multi-turn Chat
Since ChatGPT demonstrates proficiency in both MBTI-related knowledge, the multi-turn chat stage is built upon ChatGPT. The key problem PsyDI need to solve in this stage is, how to lead user to reveal underlying motivation while keep user’s interesting on our assessment, to maintain completion rate.
For the first challenge, which is to lead users to reveal their underlying motivations, we utilize multiple-choice questions where each option corresponds to a specific cognitive preference. By presenting such options, users can select the one that best aligns with their own, thereby avoiding ambiguous expressions and implicitly guiding them to reflect on their cognitive tendencies. For questions requiring confirmation, PsyDI employs forced-choice questions, compelling users to identify their most likely cognitive preferences. Additionally, PsyDI accepts free-form answers, allowing users to provide responses beyond the preset options. The prompt for this stage is list in Appendix C.
For the second challenge, which is to keep user’s interesting. We have designed a multimodal input system that allows users to input their preferred music and images, which are then converted into text to form a user statement. Additionally, our questioning framework is designed with a three-stage thought chain. First, we analyze the user statement, then we infer the user’s underlying thought patterns, and finally, we derive the next question based on these thought patterns. This approach ensures that the entire questioning process revolves around the user’s thinking style, providing greater motivation for them to continue answering questions.
@@ 4 Score Model
To achieve the goal of identifying the primary cognitive function influencing participant behavior, it’s vital to pinpoint the statement best differentiates MBTI categories. This statement will be utilized for further qualitative analysis to elucidate MBTI profiles. To address the challenge, we introduce the score model in Figure 3 to estimates MBTI representativenesss for each statement.
Within a given state with a set of statements $s=\{p_{1},p_{2},\dots,p_{n}\}$, the score model serves the purpose of evaluating the representativeness of a statement under a particular MBTI type $\mathcal{F}_{\text{MBTI}}(s)$.
Figure 3: Score model training pipeline. The pipeline begins by using ChatGPT to predict the probability of each statement’s MBTI type. We then construct pairs of statements based on these predictions and their true labels $mbti$. For each pair-wise statements $(p_{i},p_{j},mbti)$, where the statement $p_{i}$ is predicted to align more closely with the $mbti$ than statement $p_{j}$, the loss function is defined as the difference in the predicted probabilities under the $mbti$ between $p_{i}$ and $p_{j}$. This ensures that the statement more accurately matching the $label$ receives a higher score.
@@ 4.1 Ranking Training Paradigm
However, according to the existing research on MBTI and available datasets, the MBTI representativeness of a self-report statement remains an unmeasurable variable, attributable to the inherent ambiguities in language. For instance, the phrase "I easily become lost in fantasies" may align with an XNXP preference if the user emphasizes possibilities, yet shift towards an XNXJ inclination if the user focuses on long-term future considerations. To address this limitation, we propose employing the accuracy of MBTI prediction, that is, the degree to which the Large Language Model’s (LLM) predictions for a given statement correlate with the assigned labels. This accuracy serve as a proxy variable, effectively substituting for the unmeasurable statement representativeness.
@@
The rationale behind this choice lies in the quantifiability and relevance of the prediction accuracy to the statement representativeness. Notably, specific MBTI personality types may have distinctive patterns in the expression and perception of information within their statements. For example, individuals with an F preference might employ a higher frequency of emotional language, whereas those with a T preference may emphasize problem-solving solutions. By employing the prediction accuracy, we establish an observable metric connected to the statement representativeness. Therefore, we can implicitly prioritize statements that are more representative under an MBTI by ranking the statements’ prediction accuracy.
@@
It is noteworthy that, while a positive correlation exists between the prediction accuracy and the statement representativeness, it’s still challenging to establish a direct relationship between them. This difficulty arises because prediction accuracy is influenced by multiple factors, not solely the representativeness of the statement, but also the variable comprehension capabilities of Large Language Models (LLMs) in interpreting each statement. For example, a statement such as "I thrive in complex problem-solving environments" may be understood by an LLM to accurately predict an INTJ due to its analytical context. However, the same statement could also be interpreted in a way that aligns with an ENTP’s preference for innovative solutions, depending on the nuances captured by the LLM. Consequently, the score function adopts a more flexible approach by learning the ranking of the MBTI prediction accuracy rather than targeting specific accuracy values with supervised learning.
@@
The sensitivity of supervised loss to precise values is a key consideration. This loss function penalizes any deviation from the target, which poses a challenge given the ambiguous nature of the relationship between the proxy variable and the unobservable variable. In contrast, ranking loss exclusively penalizes errors in ordering, allowing the score function to remain less sensitive to noise in the data as long as it accurately captures the accuracy ranking of posts. Here "noise" refers to the inherent variability and subjectivity in self-reported statements that may not consistently align with the theoretical underpinnings of the MBTI framework. For instance, an individual might describe themselves using language that is more characteristic of one MBTI type but behave in a manner that aligns more closely with another type, creating discrepancies between self-perception and actual behavior. This variability can introduce noise into the data, complicating the direct mapping of prediction accuracy to statement representativeness. Therefore, the adoption of ranking loss can enhance the robustness of the score function.
@@ 4.2 Loss function
Therefore, considering any two statements, we can construct a pair sample $\{p_{i},p_{j},mbti\}$, where $p_{i}$ has greater representativeness of a specific MBTI type than $p_{j}$. The details of pair-wise dataset constructing process is introduced in Section 4.3. For each statement $p$, the score model is designed as an LLM with four heads, seperately predicting its score on the four dimension of MBTI (E/I, N/S, F/T, J/P). For example, the higher the score of the first head is, the statement is more likely to be like E, while a lower score represents I. Therefore, the final score of a statement is calculated as $-mbti^{\top}\cdot\sum\mathcal{F}(p_{i})$, where $mbti$ is a four-dimension vector of 1/-1, representing for the direction of true MBTI on four dimension.
Then, we construct a loss function as the difference under $mbti$ between the predicted scores of the two posts $y_{i}$ and $y_{j}$ as shown in the right side in Fig. 3. This ensures that the score for the more accurate statement is higher. To enhance the clarity of score boundaries for dynamics with varying representativeness, we introduce a margin. If the model’s score for $p_{i}$ exceeds that for $p_{j}$ by an amount less than the margin, a penalty is applied. This margin enhances the model’s robustness and generalization, mitigating the risk of overfitting. The resulting loss function is as follows:
$L(p_{i},p_{j},mbti)=\max(-mbti^{\top}\cdot(\sum\mathcal{F}(p_{i})-\sum\mathcal{F}(p_{j}))+margin)$ (6)
With Eq. 6, we can train a score model to rank the representativeness of statements under specific MBTI.
@@ 4.3 Pair-wise Dataset
For a psychology dataset composed of statements with self-reported label, it’s unreasonable to directly use label as a ground truth, since the self-reported label could be wrong. Therefore, we apply both the accuracy of MBTI prediction and the label to construct the pair-wise dataset.
Specifically, as shown in the left side in Fig. 3, for any two different statements with the same mbti label in the labeled dataset, we apply ChatGPT to predict the top-4 likely mbti and their probabilities. Then we can design an statement comparison metric to decide which statement is more likely to be the mbti, implicitly showing its representativeness of the mbti. For example, for two statements under INFP, the post that ChatGPT predicted to be 70% likely to be INFP has more representativeness comparing with the statement that ChatGPT predicted to be 50% likely to be INFP. Similarly, for one statement, we can mask some part of the statement, so that the masked statement is less likely to be the corresponding mbti under ChatGPT’s prediction. Finally, we can use the above two methods to construct pair-wise dataset.
However, due to the composition method of post mask, the score model will tend to output a higher score of a longer sentence since this tendency is aligned with the dataset, where the masked statement always has less representative than original statement. To encourage the score model to learn the true representativeness of a sentence, we apply two data augmentation methods:
Mix Datasets. In order to make the score model focus more on the representativeness of the statement than on the length of the sentence, we constructed a series of statement pairs with the same length but different representativeness. specifically, for each post, we intercepted the original statement’s 70%, and the remaining 30% is populated by random statements from other mbti to form the mix dataset.
Meaningless Datasets. To avoid the score model’s tendency to award higher scores to longer statements, we construct a dataset that is longer while representativeness remains constant, penalizing the score model’s behavior of giving higher scores to longer statements. Specifically, for each statement, we construct a sample such as $(p,p\oplus p\oplus p,mbti)$ to build a meaningless dataset by constantly repeating the statement itself.
@@ 5 Experiments
In this section, we will validate both the score model and the pipeline in PsyDI. For the score model validation, we begin in Section 5.2 by assessing the dynamic ranking accuracy of the score model compared to existing open-source and closed-source LLMs. Subsequently, we delve into the detailed evaluation of the score model’s scoring trend with specific statements. These two parts demonstrate the effectiveness of the score model from both a holistic perspective and in terms of detailed scoring metrics. In Section 5.3, we conduct an ablation study to explore the necessity of the score model’s structure and training configurations.
For the overall pipeline validation, we take a two-pronged approach. First, we examine the score variations of individual MBTI bots throughout the pipeline process to illustrate the pipeline’s effectiveness in MBTI assessment, as detailed in Section 5.4. Second, we utilize user statistics collected from PsyDI Online to observe observe whether the PsyDI pipeline shows consistent patterns in MBTI classification across different metrics, presented in Section  5.5.
5.1 Experimental Setups
Datasets. We constructed the following MBTI statement datasets with labels in both English and Chinese:
  • Reddit. The Myers Briggs Personality Tags on Reddit Data (Storey, 2018) is a public dataset consisting of user statements and their self-reported MBTI types. To balance the distribution of MBTI types, we extracted 400 statements for each MBTI type, ensuring that each statement was longer than 100 words. The Reddit dataset is augmented with the data augmentation methods introduced in section 4.3, forming Reddit-mix and Reddit-mearningless.
  • 16Personalities. We obtained all the questions from 16 Personalities  [^2] https://www.16personalities.com/, an authoritative MBTI testing site. Subsequently, we used one ChatGPT to simulate specific MBTI types and engaged another ChatGPT in multi-turn QA sessions with the former one. If the predicting ChatGPT correctly predicted the MBTI type of simulating ChatGPT, the QA record was collected along with the MBTI labels. We collect the dataset in both English and Chinese to form 16Personalities-en and 16Personalities-ch.
  • Diamente. The Diamente Chinese chit-chat dataset comprises high-quality chit-chat conversations collected with the assistance of a pretrained dialogue model that aligns well with human values. Using these chit-chats as a starting point, we asked ChatGPT to generate multi-turn QA sessions. We randomly selected answers and recorded the entire process, labeling the statements based on ChatGPT’s MBTI predictions.
Baselines. We compared the proposed score model in PsyDI with 7 existing closed-source LLMs and 4 open-source LLMs. The closed-source LLMs include GPT-3.5 Turbo (OpenAI, 2023) and GPT-4 (Achiam et al., 2023a), the acclaimed leader in LLMs, as well as DeepSeek (DeepSeek-AI, 2024), a Mixture-of-Experts (MOE) (Shazeer et al., 2017) Language Model for both English and Chinese. For open-source LLMs, we use Llama2 (Touvron et al., 2023), a commonly used English LLM, and Chatglm3 (Du et al., 2022), an open bilingual dialogue language model.
Training Setting We train PsyDI using llama-2-7b-hf as a base model and LoRA (Hu et al., 2021) finetune with LoRA (Hu et al., 2021). Since llama-2-7b-hf itself is an English large language model, we first inject English MBTI datasets from Reddit to enhance its MBTI representativeness judgment ability in English text, forming PsyDI-en. Then, we apply cross-lingual transfer learning (Schuster et al., 2018), finetuning PsyDI-en with both Reddit and Chinese datasets Diamente, so that PsyDI-ch retains its understanding of MBTI in English text while also mastering Chinese text. The lora config is setting as $r=32$, $alpha=32$, $dropout=0.05$, while the margin in loss function is set as 0.3.
Table 1: Performance of Score Model compared with both open-source models and closed-source models. Reddit-re. represents for Reddit-repeat, while 16Per. is an abbreviation for 16Personalities.
@@ 5.1 Experimental Setups
PsyDI-en 72.3% 98.1% 100% 83.4% 80.1% 70.0% 84.0%
PsyDI-zh 73.2% 98.2% 100% 85.7% 82.9% 71.9% 85.3%
Evaluating Method. To evaluate the score model in PsyDI, for each pair-wise data sample $(p_{i},p_{j},mbti)$, we asked all baseline LLMs to determine which statement was more likely from a user with the specified MBTI type $mbti$. The prompt was designed as follows:
All LLMs had their temperature set to 0.3 to ensure more deterministic outputs. For PsyDI, we calculate the score $-mbti^{\top}\cdot\sum\mathcal{F}(p)$ for each statement. Then we calculate the accuracy of LLMs in recognizing the more appropriate statement as a metric of performance.

### v3

@@ 3.2 Statement Selection
Following the update of the MBTI Profile, another key challenge emerges in eliciting more representative statement from users through subsequent questioning. Given the preliminary acquisition of user portrait statements, the next process can be divided into two distinct phases: statement selection and multi-turn dialogue with the user. The goal of the former lies in discerning which statements are most indicative of representative responses, thereby delineating the areas requiring in-depth exploration and further analysis. On the other hand, the objective of multi-turn dialogue is to engage users, encouraging them to provide more details in the extended question-answers and maintain their interest throughout the measurement.
To determine the most indicative statements, we should exclude those with minimal informational value and those that explicitly describe a specific MBTI type, as these do not warrant further inquiry. Instead, we should focus on statements that are ripe for exploration, particularly those that exhibit ambiguity across multiple MBTI dimensions. For instance, considering a user whose statements could align with either the INFP or ENFJ types. The next questioning should be designed to delve into the motivations underlying inconsistent dimensions. Thus, we design a strategy to select statements that are ambiguous yet highly representative of both the I/E and P/J. This approach aims to refine our comprehension of the nuanced meaning underlying statements and to enhance our insight into the user’s MBTI. This process can be formulated as:
$\operatorname*{arg\,min}_{p\in s}\sum_{i=1}^{4}\left(M^{i}_{j}\otimes M^{i}_{k}\right)\cdot\left(\text{rank}_{i}(\mathcal{F}_{m_{j}}(p))+\text{rank}_{i}(\mathcal{F}_{m_{k}}(p))\right)$ (5)
The notations $m_{j}$ and $m_{k}$ refer to the top-2 MBTI types within the current MBTI profile. $M$ represents the boolean (0/1) counterpart of $m$, where $M^{i}_{j}$ denotes the boolean (0/1) value of $m_{j}$ on the $i$-th MBTI dimension. The symbol $\otimes$ means XOR operation, which is employed to discern the differing dimensions between these two MBTI types. While $\text{rank}_{i}$ signifies the rank of the scores on the $i$-th MBTI dimension across all statements. The objective of Eq. 5 is to pinpoint statements that score highly on all dimensions of ambiguity, enabling subsequent questioning to elucidate the user’s true cognitive preferences.
@@ 3.3 Multi-turn Chat
Although LLMs perform exceptionally well in single-round, goal-oriented dialogues, it encounters difficulties in maintaining a coherent focus around a single objective when faced with a dozen rounds of multiple questions. To address this limitation, we have devised a hybrid approach that leverages LLMs’ strengths by structuring interactions into modules of three to five rounds, each centered on a specific goal. Within these modules, their capabilities are harnessed to progressively delve into questions related to MBTI, ensuring a step-by-step deepening of the exploration. The summarization of each module’s dialogue then contributes new insights to the pipeline, facilitating a cumulative and systematic advancement of the overall structure.
The key challenge PsyDI needs to resolve at this phase is how to guide users to reveal their underlying motivations while maintaining their interest in our measurement, thereby ensuring a high completion rate. This requires our questions to be progressively layered and logically interconnected, also engaging users throughout the process and encouraging them to provide deeper insights into their psychological profiles.
For the first challenge, to lead users to reveal their underlying motivations, we utilize multiple-choice questions where each option corresponds to a specific cognitive preference. By presenting such options, users can select the one that best aligns with their own, thereby avoiding ambiguous expressions and implicitly guiding them to reflect on their cognitive tendencies. For questions requiring confirmation, PsyDI employs forced-choice questions, compelling users to identify their most likely cognitive preferences. Also, PsyDI accepts free-form answers, allowing users to provide responses beyond the preset options. The prompt for this phase is list in Appendix C. For the second challenge, which is to keep user’s interesting. We have designed a multimodal input system that allows users to input their preferred music and images, which are then converted into text to form a user statement. Additionally, our questioning framework is designed with a three-phase thought chain. First, we analyze the user statement, then we infer the user’s underlying thought patterns, and finally, we derive the next question based on these patterns. This approach ensures that the entire questioning process revolves around the user’s thinking style, providing greater motivation for them to continue answering.
@@ 4.1 Ranking Training Paradigm
According to the existing research on MBTI (Boyle, 1995; Furnham & Crump, 2005) and available datasets, the MBTI indicativeness of a self-report statement remains an unmeasurable variable, attributable to the inherent ambiguities in language. For instance, the phrase "I easily become lost in fantasies" may align with an NP preference if the user emphasizes possibilities, yet shift towards an NJ inclination if the user focuses on long-term future considerations. To address this limitation, we propose employing the accuracy of MBTI prediction, that is, the degree to which the LLMs’ predictions for a given statement correlate with the assigned labels. This accuracy serve as a proxy variable, effectively substituting for the unmeasurable statement indicativeness. The rationale behind this choice lies in the quantifiability and relevance of the prediction accuracy to the statement indicativeness. Notably, specific MBTI personality types may have distinctive patterns in the expression and perception of information within their statements. For example, individuals with an F preference might employ a higher frequency of emotional language, whereas those with a T preference may emphasize problem-solving solutions. By employing the prediction accuracy, we establish an observable metric connected to the statement indicativeness. Therefore, we can implicitly prioritize statements that are more representative under an MBTI by ranking the statements’ prediction accuracy.
Figure 3: Score model training pipeline. It begins by using ChatGPT to predict the probability of each statement’s MBTI type. We then construct pairs of statements based on predictions and their true labels $m$. For each pair-wise statements $(p_{i},p_{j},m)$, where the statement $p_{i}$ is predicted to align more closely with the $m$ than statement $p_{j}$, the loss function is defined as the difference in the predicted probabilities under the $m$ between $p_{i}$ and $p_{j}$. This ensures that the statement more accurately matching $m$ receives a higher score.
@@
It is noteworthy that, while a positive correlation exists between the prediction accuracy and the statement indicativeness, it’s still challenging to establish a direct mapping. This difficulty arises because prediction accuracy is influenced by multiple factors, not solely the indicativeness of the statement, but also the variable comprehension capabilities of LLMs in interpreting each statement. For example, a statement such as "I thrive in complex problem-solving environments" may be understood by an LLM to accurately predict an INTJ due to its analytical context. However, the same statement could also be interpreted in a way that aligns with an ENTP’s preference for innovative solutions, depending on the nuances captured by the LLM. Consequently, the score function adopts a more flexible approach by learning the ranking of the MBTI prediction accuracy rather than targeting specific accuracy values with supervised learning.
@@
The sensitivity of supervised loss to precise values is a key consideration. This loss function penalizes any deviation from the target, which poses a challenge given the ambiguous nature of the relationship between the proxy variable and the unobservable variable. In contrast, ranking loss exclusively penalizes errors in ordering, allowing the score function to remain less sensitive to noise in the data as long as it accurately captures the accuracy ranking of statements. Here "noise" refers to the inherent variability and subjectivity in self-reported statements that may not consistently align with the theoretical underpinnings of the MBTI framework. For instance, an individual might describe themselves using language that is more characteristic of one MBTI type but behave in a manner that aligns more closely with another type, creating discrepancies between self-perception and actual behavior. This variability can introduce noise into the data, complicating the direct mapping of prediction accuracy to statement indicativeness. Therefore, the adoption of ranking loss can enhance the robustness of the score function.
@@ 4.2 Loss function
Therefore, considering any two statements, we can construct a pair sample $\{p_{i},p_{j},m\}$, where $p_{i}$ has greater indicativeness of MBTI type $m$ than $p_{j}$. The details of pair-wise dataset constructing process is introduced in Section 4.3. For each statement $p$, the score model is designed as an LLM with four heads, seperately predicting its score on the four dimension of MBTI (E/I, N/S, F/T, J/P). For example, the higher the score of the first head is, the statement is more likely to be like E, while a lower score represents I. Therefore, the final score of a statement is calculated as $-m^{\top}\cdot\sum\mathcal{F}_{m}(p_{i})$. Then, we construct a loss function as the difference under $m$ between the predicted scores of the two statements $p_{i}$ and $p_{j}$ as shown in the right side in Figure 3. This ensures that the score for the more accurate statement is higher. To enhance the clarity of score boundaries for dynamics with varying indicativeness, we introduce a margin. If the model’s score for $p_{i}$ exceeds that for $p_{j}$ by an amount less than the margin, a penalty is applied. This margin enhances the model’s robustness and generalization, mitigating the risk of overfitting. The resulting loss function is:
$L(p_{i},p_{j},m)=\max(-m^{\top}\cdot(\sum\mathcal{F}_{m}(p_{i})-\sum\mathcal{F}_{m}(p_{j}))+margin,0)$ (6)
With Eq. 6, we can train a score model to rank the indicativeness of statements under specific MBTI. The discussion of using different loss function will be presented in Section 5.3
@@ 4.3 Pair-wise Dataset
For a psychology dataset composed of statements with self-reported label, it’s unreasonable to directly use label as a ground truth since the self-reported label may not be directly linked to the text. Therefore, we apply both the accuracy of MBTI prediction and the label to construct the pair-wise dataset. Specifically, as shown in the left side in Figure 3, for any two different statements with the same MBTI label in the labeled dataset, we apply ChatGPT to predict the top-4 likely MBTI and their probabilities. Then we can design an statement comparison metric to decide which statement is more likely to be the MBTI, implicitly showing its indicativeness of the MBTI. For example, for two statements under INFP, the statement that ChatGPT predicted to be 70% likely to be INFP has more indicativeness comparing with the one predicted at a 50% likelihood. Similarly, for one statement, we can mask the last part of it, so that the masked statement is less likely to be the corresponding MBTI under ChatGPT’s prediction due to the lack of information. This approach allows us to construct extensive pair-wise datasets from any dataset annotated with MBTI labels.
However, the score model will tend to output a higher score of a longer sentence since this tendency is aligned with the dataset, where the masked statement always has less representative than original statement. To encourage the score model to learn the true indicativeness of a sentence, we apply two data augmentation methods below. The impact of these two data augmentation methods will be presented in Section 5.3.
Mix Datasets. In order to make the score model focus more on the indicativeness of the statement than on the length of the sentence, we constructed a series of statement pairs with the same length but different indicativeness. specifically, for each statement, we intercepted the original statement’s 70%, and the remaining 30% is populated by random statements from other MBTI to form the mix dataset.
Repeat Datasets. To avoid the score model’s tendency to award higher scores to longer statements, we construct a dataset that is longer while indicativeness remains constant, penalizing the score model’s behavior of giving higher scores to longer statements. Specifically, for each statement, we construct a sample such as $(p,p\oplus p\oplus p,m)$ to build a meaningless dataset by constantly repeating the statement itself.
@@ 5 Experiments
PsyDI-EN 72.3% 98.1% 98.9% 83.4% 80.1% 70.0% 84.0%
PsyDI-ZH 73.2% 98.2% 99.8% 85.7% 82.9% 71.9% 85.3%
Table 1: Comparative accuracy of Score Model against open-source and closed-source models in predicting higher scores for $p_{i}$ in sample$\{p_{i},p_{j},m\}$ with different datasets. The highest accuracy achieved in each dataset is highlighted in bold, while the second-highest accuracy is underlined.
In this section, we rigorously validate both the score model and the comprehensive pipeline within the PsyDI framework. Initially, in Section 5.2, we evaluate the ranking accuracy of the score model in comparison to both open-source and closed-source LLMs. This evaluation is complemented by a detailed examination of the model’s scoring predilection for specific statements, thereby illustrating its effectiveness from both a macroscopic and microscopic perspective. Furthermore, Section 5.3 presents an ablation study designed to ascertain the necessity of the scoring model’s architecture and training techniques.
In the validation of PsyDI pipeline, we conduct the verification from two dimensions. Initially, we observe the measurement accuracy and process of specific MBTI bots within PsyDI. This observation is followed by a validation of the stability of these measurements under random initialization, as detailed in Section 5.4. Concurrently, we examine the statistical characteristics of user data collected online, as described in Appendix D. The findings from this analysis are presented in Appendix E.2. By categorizing user data into distinct MBTI groups, each with its own statistical significance, this online version showcases the effectiveness of such classification. This two dimensions ensures a thorough and systematic validation of the PsyDI pipeline.
5.1 Experimental Setups
To fairly evaluate the performance of the PsyDI framework, we have gathered multiple datasets and compared them with both current state-of-the-art open-source and closed-source models. The datasets included in our analysis are Reddit, 16Personalities, and Diamente. Detailed information regarding the datasets, baselines, training settings, and evaluation methods can be found in the Appendix B.
Figure 4: Impact of iterative introversion augmentation on Score Model patterns. Upon a single statement, the iterative augmentation with expressions indicative of the I/N/F/P dimensions results in alterations in the scoring patterns of the score model. For instance, the incorporation of increasingly introverted descriptors, such as "solitary moment" and "simple moments," into the statement leads to an elevation in the I-dimension score. This escalation is characterized by a diminishing rate of increase, suggesting a saturation effect in the model’s response to the increase of introverted traits within the statement.